4 y Please show how to solve this equation
- 7 = -
5 10

Found 2 solutions by askme, JimMarshall:
Answer by askme(9)

4/5 * 7 = y/10
(4/5 * 7)10 = y
8*7 = y
56 = y

Answer by JimMarshall(82)

==> 4/5 * 7 = y/10
==>(4/5 * 7)10 = y
==>8*7 = y
==>56 = y